Solve:
P/3- 3P/8=5
thanks

Answer by checkley77(12844)

P/3- 3P/8=5
(8P-3*3P)/(3*8)=5
(8P-9P)/24=5
-P/24=5
-P=120
P=-120 ANS.
PROOF:
-120/3+3*120/8=5
-40+360/8=5
-40+45=5
5=5